The 6.9 percent bond of Peters Pickles has a face value of $1,000, a maturity of 15 years, semiannual interest payments, and a yield to maturity of 7.11 percent. What is the current market price of the bond?

A) $1,081.28

B) $987.95

C) $980.82

D) $990.80

E) $952.60
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =15x2
Bond Price =∑ [(6.9*1000/200)/(1 + 7.11/200)^k]     +   1000/(1 + 7.11/200)^15x2
                   k=1
Bond Price = 980.82
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %/coupons per year=1000*6.9/(2*100)
I/Y =7.11/2
N =15*2
FV =1000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(7.11/(2*100),2*15,-6.9*1000/(2*100),-1000,)